OpenStudy (anonymous):

Write the equation of the line passing through the points (0, -4) and (-2, 2).

jimthompson5910 (jim_thompson5910):

m=(y2-y1)/(x2-x1) m=(2-(-4))/(-2-0) m=(2+4)/(-2-0) m=(6)/(-2-0) m=(6)/(-2) m=-3 So the slope of the line that goes through the points is m=-3 Now let's use the point slope formula: y-y1=m(x-x1) y-(-4)=-3(x-0) y+4=-3(x-0) y+4=-3x+-3(-0) y+4=-3x+0 y+4=-3x y=-3x-4

OpenStudy (anonymous):

\[m=\frac{y-y1}{x-x1}=\frac{-4-2}{0--2}=\frac{-6}{2}\]